### v1

@@ 3.4 Hyperparameters
To optimize the performance of our model, we conducted a grid search to fine-tune hyperparameters such as the refinement layer and learning rate. Table 2 presents the range of values considered for each hyperparameter. The optimal values (in bold) are selected based on their performance during the training and validation phases.
Table 2: Hyperparameter selection (Selected hyperparameters are highlighted in bold.)
@@ 3.5 Performance Comparison and Analysis
RandomTargetDrugComb656565657070707075757575808080808585858590909090***AUROC(%)HERMESHypergraphSynergyDTFDeepSynergyARandomTargetDrugComb2525252530303030353535354040404045454545505050505555555560606060*AUPRC(%)
RandomTargetDrugComb808080808282828284848484868686868888888890909090929292929494949496969696****AUROC(%)BRandomTargetDrugComb3030303040404040505050506060606070707070***AUPRC(%)
@@
Figure 2: Model performance comparison for (A) ALMANAC Dataset and (B) ONeil Dataset. The left panels shows the AUROC (%) scores for different models across three validation modes (Random, Target, DrugComb). The right panels present the AUPRC (%) scores for the same models and validation modes. (B) ONeil Dataset. Red asterisks indicate statistical significance between HERMES and HypergraphSynergy(*** $p$-value $<0.001$; ** $p$-value $<0.01$; * $p$-value $<0.05$; two-sample $t$-test).
@@
In the NCI-ALMANAC dataset, HERMES exhibits superior performance across all evaluation strategies compared to the baseline methods (Figure 2A). Under the random partitioning strategy, HERMES achieves an AUROC of $85.91\%$, which is significantly higher than HypergraphSynergy, DTF, and DeepSynergy, which score $85.30\%$, $82.38\%$, and $83.50\%$, respectively (HERMES vs. HypergraphSynergy: $p$-value $<0.05$, two-sample $t$-test). In the more challenging ‘DrugComb’ mode, HERMES achieves an AUROC of $79.75\%$, significantly outperforming the other methods, with HypergraphSynergy recording $77.98\%$ ($p$-value $<0.01$, two-sample $t$-test), and DTF and DeepSynergy achieving even lower scores. For the ‘Target’ mode, HERMES still performs better than the rest of the methods. Additionally, the AUPRC results reflect a similar trend (right panel).
@@
The performance of HERMES is further validated using the ONEIL dataset (Figure 2B), and similar results are observed across the three modes in terms of AUROC and AUPRC. In the random stratification, HERMES attains an AUROC of 93.67%, significantly outperforming HypergraphSynergy, DTF, and DeepSynergy, which score 92.30%, 91.38%, and 90.60%, respectively (HERMES vs. HypergraphSynergy: $p$-value $<0.001$, two-sample $t$-test). In the ‘DrugComb’ mode, HERMES maintains a competitive edge with an AUROC of 88.34%, significantly surpassing HypergraphSynergy at 86.22% ($p$-value $<0.05$, two-sample $t$-test). HERMES also achieves better performance under the ‘Target’ mode compared to the other methods.
@@
To further validate the performance of our model, we considered two new validation modes using the ALMANAC dataset (which contains more samples) to test the model’s generalization ability to previously unseen drugs. The two modes are defined as follows: (1) DrugSingle where samples are stratified based on individual drugs, with one drug in each drug combination in the validation set being novel to the training set. This approach reduces the dataset size and challenges the model to predict synergies involving new drugs; (2) DrugDouble where samples are stratified such that each drug combination in the validation set contains two drugs, neither of which appears in the training set, further reducing the dataset size and increasing the difficulty. The results of ‘DrugSignle’ and ‘DrugDouble’ are shown in Figure 3. In the ‘DrugSingle’ mode, HERMES achieves an AUROC of 72.13%, which is extremely significantly higher than HypergraphSynergy’s 67.84% ($p$-value $<0.001$, two-sample $t$-test). Similarly, in the ‘DrugDouble’ mode, HERMES scores an AUROC of 68.39%, highly significantly outperforming HypergraphSynergy ($p$-value $<0.001$, two-sample $t$-test). The AUPRC metrics also show a consistent advantage for HERMES.
The comprehensive evaluation across different datasets and validation modes underscores the efficacy of HERMES in drug synergy prediction. The model consistently outperforms baseline methods in both AUROC and AUPRC metrics, demonstrating robust performance even in challenging scenarios. The results validate the superiority of HERMES in handling diverse drug synergy prediction tasks, emphasizing its importance for predicting untested drug synergies in clinical applications.
@@ 3.6 Ablation Study
The ablation study, as detailed in Table 3, scrutinizes the contributions of various components within the HERMES model utilizing the NCI-ALMANAC dataset. The evaluation encompasses three distinct modes: ‘Random’, ‘Target’, and ‘DrugComb’, with the primary metric being the average test AUROC for each configuration. The benchmark performance of the complete HERMES model registers an AUROC of $85.9\%$ in the ‘Random’ mode, $79.4\%$ in the ‘Target’ mode, and $79.8\%$ in the ‘DrugComb’ mode. These figures establish the baseline for assessing the impact of systematically removing specific components.
Table 3: Ablation study results. Average test AUROC of HERMES with and without key components.
Strategy Random Target DrugComb
HERMES 85.9% 79.4% 79.8%
w/o transformer 82.8% 78.5% 78.3%
w/o disease 85.7% 79.2% 79.5%
w/o gated residual 84.0% 79.0% 78.8%
w/o gate 82.5% 78.7% 78.5%
w/o weighted hyperedge 82.8% 78.6% 78.5%
@@
Impact of disease knowledge. The exclusion of disease knowledge demonstrates a slight yet notable impact on the model’s performance. These findings suggest that incorporating disease knowledge provides a performance advantage, but the margin is relatively narrow. Future work could explore more sophisticated methods of integrating disease knowledge to potentially amplify this performance benefit.
@@
In summary, the ablation study clearly demonstrates that each component of the HERMES model contributes uniquely to its overall efficacy. Transformer architecture, disease knowledge, gated residuals, gating mechanisms, and weighted hyperedges collectively enhance the model’s predictive performance. These results affirm the sophisticated design of HERMES, showcasing the synergistic effect of its components in achieving superior predictive accuracy across various validation strategies.
@@ 4 Conclusion
In this article, we introduced HERMES, a novel deep hypergraph learning method aimed at enhancing drug synergy prediction. By incorporating disease information and employing advanced GTNs and HGNNs with gated residual connections, HERMES achieves a superior performance on two drug synergy datasets compared to HypergraphSynergy and other methods. For future work, we will consider improving the scope of our analysis by incorporating a wider array of heterogeneous data sources such as target proteins, biological pathways, and therapeutic markers. We aim to bridge the gap between theoretical drug synergy prediction and practical clinical applications, ultimately contributing to more effective and personalized treatment strategies.

### v2

@@ 3.4 Hyperparameters
We conducted a grid search for optimized hyperparameters. Hyperparameters that yielded the highest AUROC in cross-validation were chosen for subsequent test experiments. Table 2 presents the range of values considered for each hyperparameter. The optimal values (in bold) are selected based on their performance during the training and validation phases.
Table 2: Hyperparameter selection (selected hyperparameters are highlighted in bold).
@@ 3.5 Performance Comparison and Analysis
RandomCLineDrugComb656565657070707075757575808080808585858590909090***AUROC(%)HERMESHypergraphSynergyDTFDeepSynergyNHPARandomCLineDrugComb2525252530303030353535354040404045454545505050505555555560606060*AUPRC(%)RandomCLineDrugComb30303030353535354040404045454545505050505555555560606060*****F1-score(%)
RandomCLineDrugComb808080808282828284848484868686868888888890909090929292929494949496969696****AUROC(%)BRandomCLineDrugComb3030303040404040505050506060606070707070***AUPRC(%)RandomCLineDrugComb3030303040404040505050506060606070707070*********F1-score(%)
@@
Figure 2: Model performance comparison for (A) ALMANAC Dataset and (B) ONeil Dataset. The left panels show the AUROC (%), the middle panels present the AUPRC (%), and the right panels display the F1-score (%) for different models across three validation modes (Random, CLine, DrugComb). Red asterisks indicate statistical significance between HERMES and HypergraphSynergy (*** $p$-value $<0.001$; ** $p$-value $<0.01$; * $p$-value $<0.05$; two-sample $t$-test).
@@
In the NCI-ALMANAC dataset, HERMES exhibits superior performance across all evaluation strategies compared to the baseline methods (Figure 2A). Under the random partitioning strategy, HERMES achieves an AUROC of $85.91\%$ (std: $0.0046$), which is significantly higher than HypergraphSynergy, DTF, NHP, and DeepSynergy, which score $85.30\%$, $82.38\%$, $77.69\%$, and $83.50\%$, respectively (HERMES vs. HypergraphSynergy: $p$-value $<0.05$, two-sample $t$-test). In the more challenging ‘DrugComb’ mode, HERMES achieves an AUROC of $79.75\%$ (std: $0.0112$), significantly outperforming the other methods, with HypergraphSynergy recording $77.98\%$ ($p$-value $<0.01$, two-sample $t$-test), NHP achieving $75.71\%$, and DTF and DeepSynergy achieving even lower scores. For the ‘CLine’ mode, HERMES also performs better than the rest of the methods. Additionally, the AUPRC and F1-score results reflect a similar trend (middle and right panels).
@@
The performance of HERMES is further validated using the O’NEIL dataset (Figure 2B), where similar results are observed across the three modes in terms of AUROC, AUPRC, and F1-score. In the random stratification, HERMES attains an AUROC of $93.67\%$ (std: $0.0032$), significantly outperforming HypergraphSynergy, DTF, NHP, and DeepSynergy, which score $92.30\%$, $91.38\%$, $87.07\%$, and $90.60\%$, respectively (HERMES vs. HypergraphSynergy: $p$-value $<0.001$, two-sample $t$-test). In the DrugComb mode, HERMES maintains a competitive edge with an AUROC of $88.34\%$ (std: $0.0205$), significantly surpassing HypergraphSynergy at $86.22\%$ ($p$-value $<0.05$, two-sample $t$-test) and outperforming NHP at $82.92\%$. HERMES also achieves better performance under the CLine mode compared to the other methods, further supporting its robustness across different datasets.
To further validate the performance of our model, we considered two new validation modes using the ALMANAC dataset (Figure 3) to assess the model’s generalization ability to previously unseen drugs. The two modes are defined as follows: (1) DrugSingle, where samples are stratified based on individual drugs, with one drug in each drug combination in the validation set being novel to the training set. This approach reduces the dataset size and challenges the model to predict synergies involving new drugs; (2) DrugDouble, where samples are stratified such that each drug combination in the validation set contains two drugs, neither of which appears in the training set, further reducing the dataset size and increasing the difficulty. The results of ‘DrugSingle’ and ‘DrugDouble’ are shown in Figure 3. In the ‘DrugSingle’ mode, HERMES achieves an AUROC of $72.13\%$ (std: $0.0118$), which is significantly higher than HypergraphSynergy’s $67.84\%$ ($p$-value $<0.001$, two-sample $t$-test). Similarly, in the ‘DrugDouble’ mode, HERMES scores an AUROC of $68.39\%$ (std: $0.0468$), significantly outperforming HypergraphSynergy ($p$-value $<0.001$, two-sample $t$-test). The AUPRC metrics also show a consistent advantage for HERMES.
The comprehensive evaluation across different datasets and validation modes underscores the efficacy of HERMES in drug synergy prediction. Although statistical insignificance is observed in certain validation modes, indicating that the extent of improvement can vary based on data distribution, sample size, and task difficulty, HERMES consistently demonstrates competitive and robust performance compared to other methods. These results validate the superiority of HERMES in addressing diverse drug synergy prediction tasks and emphasize its potential for predicting untested drug synergies in clinical applications.
@@ 3.6 Ablation Study
The ablation study, as detailed in Table 3, scrutinizes the contributions of various components within the HERMES model utilizing the NCI-ALMANAC dataset. The evaluation encompasses three distinct modes: ‘Random’, ‘CLine’, and ‘DrugComb’, with the primary metric being the average test AUROC for each configuration. The benchmark performance of the complete HERMES model registers an AUROC of $85.9\%$ in the ‘Random’ mode, $79.4\%$ in the ‘CLine’ mode, and $79.8\%$ in the ‘DrugComb’ mode. These figures provide a baseline for evaluating the effects of systematically removing specific components or combinations of components.
Table 3: Ablation study results. Average test AUROC (with standard derivation) of HERMES with and without key components.
Strategy Random CLine DrugComb
[…]
w/o transformer $82.8\%\pm 0.007$ $78.5\%\pm 0.013$ $78.3\%\pm 0.011$
w/o disease $85.7\%\pm 0.006$ $79.2\%\pm 0.016$ $79.5\%\pm 0.007$
w/o gated residual $84.0\%\pm 0.060$ $79.0\%\pm 0.014$ $78.8\%\pm 0.009$
w/o gate $82.5\%\pm 0.006$ $78.7\%\pm 0.013$ $78.5\%\pm 0.009$
w/o transformer + w/o disease $82.1\%\pm 0.005$ $77.8\%\pm 0.015$ $77.3\%\pm 0.021$
w/o transformer + w/o gate $82.0\%\pm 0.004$ $77.7\%\pm 0.014$ $77.2\%\pm 0.021$
w/o transformer + w/o disease + w/o gate $80.4\%\pm 0.006$ $76.5\%\pm 0.016$ $77.8\%\pm 0.024$
@@
Impact of disease knowledge. The exclusion of disease (drug-indication relation) demonstrates a slight yet notable impact on the model’s performance. These findings suggest that incorporating disease knowledge provides a performance advantage, but the margin remains relatively narrow. This may stem from the single-modal nature of the current disease data and the limitations of the classical CODER model. Exploring the integration of multimodal disease knowledge and adapting more advanced large language models could potentially enhance this performance benefit.
@@
In summary, the ablation study clearly demonstrates that each component of the HERMES model contributes uniquely to its overall efficacy. Transformer architecture, molecule indication relation, gated residuals, gating mechanisms, and weighted hyperedges collectively enhance the model’s predictive performance. These results affirm the sophisticated design of HERMES, showcasing the synergistic effect of its components in achieving superior predictive accuracy across various validation strategies.
@@ 4 Conclusion
In this article, we introduced HERMES, a novel deep hypergraph learning framework for drug synergy prediction that integrates heterogeneous biomedical data, including drug molecular structures, gene expression profiles of cell lines, and disease indications. Our results demonstrate that HERMES consistently achieves superior performance across two large-scale benchmark datasets, excelling particularly in challenging scenarios with previously untested drug combinations. The model’s gated residual mechanism mitigates over-smoothing in message-passing, enabling it to capture intricate high-order relationships among drugs, cell lines, and diseases with enhanced precision.
While HERMES establishes a flexible and general framework for drug synergy prediction, there are areas for improvement. First, the model currently incorporates only drug, cell line, and disease information, limiting its ability to leverage other relevant biomedical knowledge, such as protein–protein interactions, drug–target interactions, and disease–gene interactions. Integrating such information would further enrich the hypergraph structure and improve the model’s generalizability and predictive accuracy. Additionally, HERMES’s computational demands, particularly in terms of CPU and GPU usage, could present challenges when scaling to extremely large datasets or incorporating more data sources. To address this, innovative graph/hypergraph sampling techniques and optimized memory management strategies should be explored. Finally, HERMES presently models interactions between only two drugs. Extending the framework to accommodate multi-drug combinations could open promising new avenues for exploring complex treatment regimens and identifying synergies in multi-drug therapies.
In summary, HERMES presents a scalable and adaptable approach to drug synergy prediction, with potential applications extending into other biomedical and clinical research domains, such as drug discovery, personalized medicine, and the development of more effective and efficient therapeutic strategies. It enhances the ease of laboratory implementation while also providing cost-effective solutions for assessing drug synergy. The framework’s flexibility opens further opportunities to incorporate additional data sources and adapt to diverse predictive tasks, making it a foundational step towards more accurate and generalizable drug synergy models.